# ACT-R: Adaptive Camera Trajectories for Single-View 3D Reconstruction

### v2

@@ Abstract
We introduce the simple idea of adaptive view planning to multi-view synthesis, aiming to improve both occlusion revelation and 3D consistency for single-view 3D reconstruction. Instead of producing an unordered set of views independently or simultaneously, we generate a sequence of views, leveraging temporal consistency to enhance 3D coherence. More importantly, our view sequence is not determined by a pre-determined and fixed camera setup. Instead, we compute an adaptive camera trajectory (ACT), forming an orbit, which seeks to maximize the visibility of occluded regions of the 3D object to be reconstructed. Once the best orbit is found, we feed it to a video diffusion model to generate novel views around the orbit, which can then be passed to any multi-view 3D reconstruction model to obtain the final result. Our multi-view synthesis pipeline is quite efficient since it involves no run-time training/optimization, only forward inferences by applying pre-trained models for occlusion analysis and multi-view synthesis. Our method predicts camera trajectories that reveal occlusions effectively and produce consistent novel views, significantly improving 3D reconstruction over SOTA alternatives on the unseen GSO dataset.
Figure 1: Our single-view 3D reconstruction method, ACT-R, predicts an adaptive camera trajectory (green), aiming to maximize the visibility of occluded object parts over a fixed sequence length (20 views in this example). The trajectory, obtained in under 10s, is then used by a video generator (e.g., SV3D [1]) to produce a sequence of novel views for multi-view 3D reconstruction, here using the established NeUS model [2]. Compared to a generic trajectory (red), at fixed elevation, ACT-R yields much cleaner results with more faithful recovery of occluded regions.
@@ 1 Introduction
Figure 2: Pipeline of our single-view reconstruction method, ACT-R, with adaptive camera trajectories (ACT). We first employ Slice3D [3] to produce the slice images of the input object, with the slicing direction from the camera to the object center. Then we compute the semantic difference between the input and its slices by comparing their $512\times 7\times 7$ feature maps extracted from VGG16 [4]. For better visualization, each difference map $d_{i}\in[0,1]^{7\times 7}$ is up-scaled and overlaid onto slice images . Afterwards, we identify the regions that have significant semantic differences (Sec. 3.2), and plan the camera trajectories based on them (Sec. 3.3). Finally, we condition SV3D [1] on our planned trajectories, yielding a sequence of views, which can be fed into either NeUS [2] or InstantMesh (IM) [5] for multi-view 3D reconstruction.
3D reconstruction from single-view images has been one of the most intensively studied problems in computer graphics. One class of modern approaches directly generate or regress 3D representations of objects from input images [6, 7, 8, 9, 10, 11], often resorting to 3D supervision which requires large 3D datasets for training. With significant advances in novel view synthesis [12, 13], the second line of popular approaches to single-view 3D reconstruction first perform a multi-view synthesis, which typically generates an unordered set of views either independently [14] or simultaneously [15, 16, 17, 18] with fixed camera setups. This is followed by multi-view 3D reconstruction via differentiable rendering (e.g., NeUS [2] and variants [19, 20]) so that the entire solution pipeline can avoid direct 3D supervisionWe use the term “direct 3D supervision” in the context of 3D reconstruction to refer to models trained using paired images and ground-truth 3D models..
The main challenges to single-to-multi view synthesis are twofold. First, the multi-view images should reveal the occluded structures of the target 3D object that are hidden from the input image, which is a highly ill-posed problem. Second, the generated views must attain 3D consistency to ensure that a plausible and coherent 3D model can be reconstructed. Upon close examination of state-of-the-art multi-view synthesis methods, as well as single-view reconstruction methods which combine such syntheses with direct 3D prediction [21, 22, 23, 24], we find that there is still much room for improvement.
@@
Extensive experiments demonstrate that our method predicts camera trajectories tailored to each input example, effectively revealing occluded regions for higher-quality reconstruction. On the GSO benchmark [28], ACT-R outperforms, both qualitatively and quantitatively, state-of-the-art methods based on multi-view syntheses from static [1], fixed [23, 16], or randomly perturbed [1] camera trajectories, all without direct 3D supervision, as well as a recent method, Craftsman [29], which employs 3D supervision in the multi-view reconstruction step.
Figure 3: Initial single-view inputs to advanced methods such as Trellis and Hunyuan3D (HY) can be improved by one of the synthesized multi-view images from ACT-R, to improve reconstruction results; see highlights.
On the other hand, ACT-R, which is best positioned to boost multi-view 3D reconstruction methods without direct 3D supervision, cannot beat some of the latest LRMs such as CLAY [6], Hunyuan3D [10], or Trellis [9], which all use direct 3D supervision trained by closed-source (CLAY) or carefully curated (Trellis) large-scale 3D assets. Indeed, directly comparing ACT-R with these models can be unfair due to the different training setups. Nevertheless, as shown in Fig. 3, our ACT can benefit these supervised LRMs by improving the input views.
@@ 2 Related Work
As 3D reconstruction, even only for single views, has been extensively studied, we will only focus on closely related works to ACT-R, on video generation, multi-view synthesis, and view planning.
@@ 2.1 Video Generation in 2D and for 3D
Early works [30, 31, 32] extend image diffusion to video generation. Stable Video Diffusion (SVD) [33] adapts latent diffusion methods [34, 35] to large-scale video datasets for temporally coherent generation. A stream of work focuses on keyframe conditioning [36, 37, 38, 39], where initial frames are generated to anchor subsequent video synthesis, with latent-consistency networks ensuring temporal and appearance coherence. PYoCo [40] employs tailored noise priors for fine-tuning image models, while Lumiere [41] represents videos in a single latent space without keyframe dependencies. Methods like Tune-a-video [42] enable few-shot video synthesis through fine-tuning of pre-trained image diffusion models. Training-free approaches [43, 44] utilize Large Lanague Models (LLMs) for generation guidance, while other works [45, 46] have explored various latent diffusion strategies for video synthesis. Although video generation models typically lack explicit 3D representations, they can still achieve 3D consistency by producing temporally coherent videos. In our approach, we use video diffusion models as the backbone to generate view sequences.
@@ 2.3 View and Path Planning
Path planning has applications in navigation, scanning, and even computational fabrication [54, 55, 56, 57, 58, 59]. In particular, Next-Best-View (NBV) planning addresses the fundamental challenge of determining an optimal sequence of camera positions to maximize information gain during scene or object inspection  [60, 61, 62]. Two popular traditional approaches to this problem include voxel-space methods that optimize coverage metrics  [60, 61, 62, 63], and surface-based methods that analyze boundary characteristics to determine optimal viewpoints  [64, 65, 66]. Recent advances in deep learning have transformed the NBV paradigm. Daryl et al. [67] leveraged reinforcement learning to predict optimal viewpoints by maximizing coverage-based rewards. Liu et al. [58] developed a learning-based reconstructability predictor to guide their iterative view planner for 3D reconstruction. NeU-NBV [68] introduced a novel uncertainty evaluation framework that utilizes neural image rendering to inform NBV prediction. GenNBV [69] proposed a generalizable NBV policy to enhance its performance on unseen scenes.
While traditional NBVs assume the availability of a complete 3D reference model, our work addresses a more challenging scenario where only a single image serves as input. This constraint fundamentally shifts the problem from pure coverage optimization to view prediction based on limited initial information, requiring novel strategies for viewpoint selection and trajectory planning.
@@ 3.2 Building Semantic Difference Blocks
[…]
@@
Compared with more recent feature encoders such as DINOv2 [70], we empirically find that VGG16 [4] works better in extracting semantic difference maps between slices and input images. Please see the supplementary materials for detailed ablation studies.
@@ 3.3 Camera Trajectory Planning
Since the search space for camera trajectories is infinite, we discretize the camera’s movement to facilitate the search. Specifically, we begin by sampling the azimuth steps at a constant interval of $18^{\circ}$ (calculated as $20360^{\circ}/$ 20) between each frame from $0^{\circ}$ to $360^{\circ}$, which creates a closed-loop trajectory in terms of azimuth.
Next, we define the elevation angle changes at each step using the set {±$\{\pm 5^{\circ},\pm 4^{\circ},\pm 3^{\circ},\pm 2^{\circ},\pm 1^{\circ},0^{\circ}\}$. We limit per-step elevation changes to within $5^{\circ}$, as larger step sizes challenges frame-consistent video generation. Each orbit is divided into four segments based on azimuth angles. Within each segment, the elevation angle increments at a constant rate, with the step size selected from the set. To ensure a closed orbit, we enforce the total variation in elevation to be zero by mirroring and negating the elevation change in the second segment onto the third and the first onto the fourth. This approach results in a total of $11\times 11=121$ candidate trajectories, which we denote as the set $\Pi$.
@@ 3.4 View Generation and 3D Reconstruction
We fed the camera trajectory $\pi^{*}$ to SV3D [1] to generate a video containing a sequence of novel views. Since the video could suffer from significant artifacts due to stochasticity in the generative models, we apply view-consistency as the primary metric to filter out low-quality results and regenerate with alternative random seeds when necessary. The final view sequence can be integrated into any 3D reconstruction pipeline that accepts posed-multiview-images as input. We demonstrate ACT-R’s flexibility by reconstruction mesh through two different approaches: volumetric rendering with NeUS [2] using all 21 generated images, and processing through LRM [5] by feeding 6 uniformly sampled key frames from the 21 views.
@@ 4.1 Qualitative Visual Comparisons
[…]
  3. 3.
    $\text{Ours}_{\text{IM}}$: LRM with 6 views sampled uniformly from adaptive camera trajectory produced by our method.
Qualitative results in  Figures 9 and 10 show that different reconstruction backbones offer distinct advantages, while also expose their limitations. NeUS produces more faithful reconstructions that closely match the input, yet suffers from blobby surface artifacts. Effectively leveraging 3D priors, LRMs produce smoother surfaces, but respect less the details from the input image. Despite these inherent backbone-specific characteristics, ACT-R’s adaptive trajectories consistently improve reconstruction quality, particularly over occluded regions. Our approach more faithfully reconstructs the solid bowl interiors (rows 2 and 3) that most competing methods overlooked due to occlusion in the input image. It also provides enhanced spatial awareness between objects (row 5) and creates reasonable joint connections in complex structures such as the umbrellas (row 6).
@@ 4.3 Visualizations of Trajectories
For the Bundt cake pan, even with a positive elevation for the input view, we still raise the camera to reveal the inner tube. Although slice3D [3] failed to predict the inner tube in the sliced image, it still suggests that inner regions require additional attention. For the tape example, we can see the camera highest elevation is lower that the pan example, since it has already fully observed the hole in the middle. This proves that our method is adaptive to each single object. For the sneaker example, our trajectory can better observe the pair of wings so that they will not occlude with each other in most of the frames. For the sofa, we lower the camera to fully capture its bottom. For the Jenga blocks, we planned a trajectory that can better observe the concave areas.
@@ 4.4 Ablation Studies
Figure 6: Computing semantic difference maps $\{d_{i}\}$ using different metrics.
@@
Robustness to camera pose estimation. Since our trajectory planning operates on a coarse 3D representation, i.e., the blocks, it is robust against errors from estimating the camera parameters and the bounding boxes. More details on how camera estimation affects our view planning can be found in the Supplementary.
@@ 5 Conclusion, Limitation, and Future Work
In future work, a more in-depth investigation into how to best integrate adaptive view planning into direct image-to-3D LRMs is worth conducting. We would also like to explore other applications of ACT, e.g., for robotics and autoscanning [79].
Figure 10: Additional qualitative visual comparisons between single-view 3D reconstruction methods on the GSO dataset. Please zoom in for a closer inspection. Meshes reconstructed with NeUS and LRM are in yellow and blue blocks, respectively. Our slice-guided approach demonstrates superior reconstruction of occluded regions, including bowl bottoms, hidden wings and extended features like raccoon tails and elephant trunk. Please zoom in for detailed comparison of the reconstruction quality.
@@
Appendix A Estimation of Camera Poses and Object Bounding Boxes
@@
Appendix B Robustness Against Camera Pose Errors
@@ Appendix C Visibility Set Calculation
To calculate the visibility set $\psi(\pi(t))$ in Equation (1) from the main paper, we perform two key checks for each difference block: (1) a field of view check and (2) an occlusion check. The field of view check ensures blocks lie within the camera’s viewing angle (33.6°), computed as the angle between the camera direction and the vector from camera to block center. The occlusion check determines if any other block intersects the line segment from camera to the target block’s center using a ray-box intersection algorithm. A block is considered visible only if it passes both checks. This visibility function allows us to evaluate each candidate trajectory by summing the weights of all visible blocks across all camera positions, leading to the selection of the optimal trajectory $\pi^{*}$ that maximizes the visibility of important difference regions.
Appendix D Calculation of Random Trajectory

### v3

@@ Abstract
We introduce the simple idea of adaptive view planning to multi-view synthesis, aiming to improve both occlusion revelation and 3D consistency for single-view 3D reconstruction. Instead of producing an unordered set of views independently or simultaneously, we generate a sequence of views, leveraging temporal consistency to enhance 3D coherence. Importantly, our view sequence is not determined by a pre-determined and fixed camera setup. Instead, we compute an adaptive camera trajectory (ACT), to maximize the visibility of occluded regions of the 3D object to be reconstructed. Once the best orbit is found, we feed it to a video diffusion model to generate novel views around the orbit, which can then be passed to any multi-view 3D reconstruction model to obtain the final result. Our multi-view synthesis pipeline is quite efficient since it involves no run-time training/optimization, only forward inferences by applying pre-trained models for occlusion analysis and multi-view synthesis. Our method predicts camera trajectories that reveal occlusions effectively and produce consistent novel views, significantly improving 3D reconstruction over SOTA alternatives on the unseen GSO dataset.
Figure 1: ACT-R, for single-view 3D reconstruction, predicts an adaptive camera trajectory (green) to maximize the visibility of occluded object parts over a fixed sequence length (20 views). The trajectory, obtained in under 10s, is then used by a video generator (e.g., SV3D [60]) to produce a sequence of novel views for multi-view 3D reconstruction, here using NeUS [62]. Compared to a generic trajectory (red), at fixed elevation, ACT-R yields much cleaner results with more faithful recovery of occluded regions.
@@ 1 Introduction
Figure 2: Pipeline of our single-view reconstruction method, ACT-R, with adaptive camera trajectories (ACT). We first employ Slice3D [64] to produce the slice images of the input object, with the slicing direction from the camera to the object center. Then we compute the semantic difference between the input and its slices by comparing their $512\times 7\times 7$ feature maps extracted from VGG16 [52]. Each difference map $d_{i}\in[0,1]^{7\times 7}$ is up-scaled and overlaid onto slice images for beter visualization. Next, we identify the regions that have significant semantic differences (Sec. 3.2), and plan the camera trajectories based on them (Sec. 3.3). Finally, we condition SV3D [60] on our planned trajectories, yielding a sequence of views, which can be fed into NeUS [62] or InstantMesh (IM) [73] for multi-view 3D reconstruction.
Single-view 3D reconstruction has been one of the most intensively studied problems in computer vision. One class of modern approaches directly generate or regress 3D representations of objects from input images [77, 29, 70, 71, 58, 80], often resorting to 3D supervision which requires large 3D datasets for training. With significant advances in novel view synthesis [41, 23], the second line of popular approaches to single-view 3D reconstruction first perform a multi-view synthesis, which typically generates an unordered set of views either independently [32] or simultaneously [34, 36, 56, 51] with fixed camera setups. This is followed by multi-view 3D reconstruction via differentiable rendering (e.g., NeUS [62] and variants [35, 61]) so that the entire solution pipeline can avoid direct 3D supervisionWe use the term “direct 3D supervision” in the context of 3D reconstruction to refer to models trained using paired images and ground-truth 3D models..
The main challenges to multi-view synthesis are twofold. First, the produced images should reveal occluded structures of the target 3D object that are hidden from the input image. Second, the generated views must attain 3D consistency to ensure that a plausible and coherent 3D model can be reconstructed. Upon close examination of state-of-the-art multi-view synthesis methods, as well as single-view reconstruction methods which combine such syntheses with direct 3D prediction [25, 66, 54, 31], we find that there is still much room for improvement.
@@
Figure 3: Initial single-view inputs to advanced methods such as Trellis and Hunyuan3D (HY) can be improved by one of the synthesized multi-view images from ACT-R, to improve reconstruction results, as highlighted in the colored regions.
Extensive experiments demonstrate that our method predicts camera trajectories tailored to each input example, effectively revealing occluded regions for higher-quality reconstruction. On the GSO benchmark [9], ACT-R outperforms, both qualitatively and quantitatively, state-of-the-art methods based on multi-view syntheses from static [60], fixed [54, 36], or randomly perturbed [60] camera trajectories, all without direct 3D supervision, as well as a recent method, Craftsman [26], which employs 3D supervision in the multi-view reconstruction step. On the other hand, ACT, which is best positioned to boost multi-view 3D reconstruction without direct 3D supervision, cannot beat some of the latest LRMs such as CLAY [77], Hunyuan3D [58], or Trellis [71], which all use direct 3D supervision trained by closed-source (CLAY) or carefully curated (Trellis) large-scale 3D assets. Directly comparing ACT-R with these models can be unfair due to the different training setups. Nevertheless, as shown in Fig. 3, ACT can benefit these supervised LRMs by improving the input views.
@@ 2 Related Work
As single-view 3D reconstruction has been extensively studied, we only focus on closely related works on video generation, multi-view synthesis, and view planning.
@@ 2.1 Video Generation in 2D and for 3D
Early works [17, 53, 16, 11, 2, 68, 79, 81] extend image diffusion to video generation. Stable Video Diffusion (SVD) [3] adapts latent diffusion methods [49, 4] to large-scale video datasets for temporally coherent generation. A stream of work focuses on keyframe conditioning [12, 63, 27, 76], where initial frames are generated to anchor subsequent video synthesis, with latent-consistency networks ensuring temporal and appearance coherence. Training-free approaches [18, 20] utilize Large Lanague Models (LLMs) for generation guidance. Although video generation models typically lack explicit 3D representations, they can still achieve 3D consistency by producing temporally coherent videos. In our approach, we use video diffusion models as the backbone to generate view sequences.
@@ 2.3 View and Path Planning
Path planning has applications in navigation, scanning, and even computational fabrication [10, 22, 47, 48, 33, 37]. In particular, Next-Best-View (NBV) planning addresses the fundamental challenge of determining an optimal sequence of camera positions to maximize information gain during scene or object inspection  [8, 39, 59]. Two popular traditional approaches to this problem include voxel-space methods that optimize coverage metrics  [8, 39, 59, 38], and surface-based methods that analyze boundary characteristics to determine optimal viewpoints  [44, 24, 6]. Recent advances in deep learning have transformed the NBV paradigm, introducing reinforcement learning [43, 7], reconstructability predictor [33] and uncertainty evaluation framework [21] to the scope. While traditional NBVs assume the availability of a complete 3D reference model, our work addresses a more challenging scenario where only a single image serves as input. This constraint fundamentally shifts the problem from pure coverage optimization to view prediction based on limited initial information, requiring novel strategies for trajectory planning.
@@ 3.2 Building Semantic Difference Blocks
As illustrated in the first step in Fig. 2, Slice3D produces in total $M$ slice images $\{s_{1},s_{2},\dots,s_{M}\}$ from the input $x$. We quantify the semantic differences between slice image $s_{i}$ and input $x$ by comparing the features from the final pooling layer of VGG16 [52]. Specifically, we compute: $d_{i}=\langle\phi(x),\phi(s_{i})\rangle,$ where $\phi(\cdot)\in\mathbb{R}^{512\times 7\times 7}$ represents the VGG16 feature maps for a given input, with spatial resolution $7\times 7$ and feature dimension $512$ per position; $1\leq i\leq M$; $\langle\cdot,\cdot\rangle$ denotes the cosine similarity between the 512-dimensional feature vectors at each corresponding spatial position from the two inputs, resulting in $d_{i}\in\mathbb{R}^{7\times 7}$.
@@
Compared with more recent feature encoders such as DINOv2 [42] and CLIP [46], we empirically find that VGG16 [52] works better in extracting semantic difference maps between slices and input images. Please see the supplementary materials for detailed ablation studies.
@@ 3.3 Camera Trajectory Planning
Since the search space for camera trajectories is infinite, we discretize the camera’s movement to facilitate the search. Specifically, we begin by sampling the azimuth steps at a constant interval of $18^{\circ}$ (calculated as $360^{\circ}/20$) between each frame from $0^{\circ}$ to $360^{\circ}$, which creates a closed-loop trajectory in terms of azimuth.
Next, we define the elevation angle changes at each step using the set $\{\pm 5^{\circ},\pm 4^{\circ},\pm 3^{\circ},\pm 2^{\circ},\pm 1^{\circ},0^{\circ}\}$. We limit per-step elevation changes to within $5^{\circ}$, as larger step sizes challenges frame-consistent video generation. Each orbit is divided into four segments based on azimuth angles. Within each segment, the elevation angle increments at a constant rate, with the step size selected from the set. To ensure a closed orbit, we enforce the total variation in elevation to be zero by mirroring and negating the elevation change in the second segment onto the third and the first onto the fourth. This approach results in a total of $11\times 11=121$ candidate trajectories, which we denote as the set $\Pi$.
@@ 3.4 View Generation and 3D Reconstruction
We fed the camera trajectory $\pi^{*}$ to SV3D [60] to generate a video containing a sequence of novel views. Since the video could suffer from significant artifacts due to stochasticity in the generative models, we apply view-consistency as the primary metric to filter out low-quality results and regenerate with alternative random seeds when necessary. The final view sequence can be integrated into any 3D reconstruction pipeline that accepts posed-multiview-images as input. We demonstrate ACT-R’s flexibility by reconstruction meshes through two different approaches: volumetric rendering with NeUS [62] using all 21 generated images, and processing through LRM [73] by feeding 6 uniformly sampled key frames from the 21 views.
@@ 4.1 Qualitative Visual Comparisons
    $\text{SV3D}_{\text{IM}}$: LRM with 6 views sampled uniformly from generic camera trajectories with constant elevation.
Qualitative results in  Figure 6 show that different reconstruction backbones offer distinct advantages, while also expose their limitations. NeUS produces more faithful reconstructions that closely match the input, yet suffers from blobby surface artifacts. Effectively leveraging 3D priors, LRMs produce smoother surfaces, but respect less the details from the input image. Despite these inherent backbone-specific characteristics, ACT-R’s adaptive trajectories consistently improve reconstruction quality, particularly over occluded regions. Our approach more faithfully reconstructs the solid bowl interiors (rows 2) that most competing methods overlooked due to occlusion in the input image and provides enhanced spatial awareness between objects (row 3). See supplementary material for more results.
@@ 4.2 Quantitative Comparisons
We further compared the coverage metric resulting from different trajectories. Quantitative numbers suggest that our adaptive trajectory achieves a better coverage than static or random trajectories, please see the supplementary material for more details.
@@ 4.3 Visualizations of Trajectories
For the Bundt cake pan, even with a positive elevation for the input view, we still raise the camera to reveal the inner tube. Although slice3D [64] failed to predict the inner tube in the sliced image, it still suggests that inner regions require additional attention. For the tape example, the camera highest elevation is lower that the pan example, since it has already fully observed the hole in the middle. This proves that our method is adaptive to each single object. For the sneaker example, our trajectory can better observe the pair of wings so that they will not occlude with each other in most of the frames. For the sofa, we lower the camera to fully capture its bottom. For the Jenga blocks, we planned a trajectory that can better observe the concave areas.
@@ 4.4 Ablation Studies
Figure 7: Computing semantic difference maps $\{d_{i}\}$ using different metrics. Brighter color indicates higher semantic difference.
@@
Robustness to camera pose estimation. Since our trajectory planning operates on a coarse 3D representation, i.e., the blocks, it is robust against errors from estimating the camera parameters and the bounding boxes. More details on how camera estimation affects our view planning can be found in the supplementary material.
@@ 5 Conclusion, Limitation, and Future Work
In future work, a more in-depth investigation into how to best integrate adaptive view planning into direct image-to-3D large reconstruction models (LRMs) is worth conducting. We would also like to explore other applications of ACT, e.g., for robotics and autoscanning [69].
@@
6 Additional Results
We provide additional qualitative results from the GSO dataset in Figures 14 and 15.
7 Estimation of Camera Poses and Object Bounding Boxes
@@
8 Robustness Against Camera Pose Errors
@@ 9 Visibility Set Calculation
To calculate the visibility set $\psi(\pi(t))$ in Equation (1) from the main paper, we perform two key checks for each difference block: (1) a field of view check and (2) an occlusion check. The field of view check ensures blocks lie within the camera’s viewing angle (33.6°), computed as the angle between the camera direction and the vector from camera to block center. The occlusion check determines if any other block intersects the line segment from the camera to the target block’s center using a ray-box intersection algorithm. A block is considered visible only if it passes both checks. This visibility function allows us to evaluate each candidate trajectory by summing the weights of all visible blocks across all camera positions, leading to the selection of the optimal trajectory $\pi^{*}$ that maximizes the visibility of important difference regions.
10 Calculation of Random Trajectory